OpenStudy (anonymous):

How do you solve this rational equation: 1/10 + 1/15 = 1/x

Parth (parthkohli):

Multiply both sides by \(\rm 150 x\).

OpenStudy (anonymous):

you have to simplfy the denominators. for example, what is a number that goes into both 10 and 15?

OpenStudy (anonymous):

so you would divide the number that goes into 10 and 15 and the result you will multiply and get 1/result from multiplying simplified denominators.

Parth (parthkohli):

You have to clear the denominators, actually.

OpenStudy (anonymous):

for example. 1/24 + 1/6 = 1/x 6 goes into both. so divide denominators by 6 so you would get 1/4 + 1/1(technically 1) and 1*4 = 5 so result would be 1/5

OpenStudy (anonymous):

yes the long way to do it would be to multiply by 150 but thats as useful in this case. other cases, yes, that would be necessary.

OpenStudy (anonymous):

okay, makes sense. Thank you! I was doing the opposite at first - I multiplied each denominator so they would be the same. But, I didn't think that was right.

OpenStudy (anonymous):

sweet:) glad i could help!